# SemEval-2025 Task 11: Bridging the Gap in Text-Based Emotion Detection

### v1

@@ 6 Discussion
Many participants also experimented with traditional transformer-based models, particularly XLM-RoBERTa, mBERT, DeBERTa, and IndicBERT (see Figure 3 and Figure 4 in the Appendix).
@@
The results from the top-performing submissions suggest that while LLMs achieve strong overall performance, their effectiveness heavily depends on prompt engineering techniques and wording.
Additionally, performance varies significantly by language. Across all tracks, LLM-based approaches and the best-performing systems consistently yielded better results for high-resource languages such as English and Russian. Conversely, performance dropped notably when tested on low-resource languages such as Swahili and Emakhuwa.
Furthermore, most teams did not incorporate additional datasets to enhance performance (see Appendix) as few-shot and zero-shot approaches proved highly effective.
@@ 7 Conclusion
We presented our shared task on text-based emotion recognition, which covered three tracks and a total of 32 languages. The submitted systems were ranked based on macro F1-scores for Tracks A and C, comparing predicted labels to gold labels, and based on the ranking of predicted intensity scores for Track B. We summarised the reported results, highlighting the best-performing and most innovative methods. Overall, performance varied significantly across languages.
Limitations
Emotions are subjective and subtle, and they are expressed and perceived differently. We do not claim that our datasets capture the true emotions of the speakers, fully represent language use across the 32 languages, or cover all possible emotions. We discuss these limitations extensively in Ethical Considerations.

### v2

@@ 6 Discussion
Many participants also experimented with traditional transformer-based models, particularly XLM-RoBERTa, mBERT, DeBERTa, and IndicBERT Kakwani et al. (2020) (see Figure 3 and Figure 4 in the Appendix).
@@
The results from the top-performing submissions suggest that while LLMs achieve strong overall performance, their effectiveness is heavily dependent on prompt engineering techniques and wording.
Additionally, performance varies significantly by language. Across all tracks, LLM-based approaches and the best-performing systems consistently yielded better results for high-resource languages such as English and Russian. In contrast, performance dropped notably when tested on low-resource languages such as Swahili and Emakhuwa.
Furthermore, most teams did not incorporate additional datasets to enhance performance (see Appendix), as few-shot and zero-shot approaches proved highly effective.
@@ 7 Conclusion
We presented our shared task on text-based emotion recognition, which covered three tracks and a total of 32 languages. The submitted systems were ranked based on macro F1-scores for Tracks A and C, comparing predicted labels to gold labels, and based on the ranking of predicted intensity scores for Track B.
We summarised the reported results, discussing the best-performing and most innovative methods.
Overall, performance varied significantly across languages. Our results highlight that emotion recognition remains an open challenge, particularly for under-served languages and in low-resource settings.
8 Limitations
Emotions are subjective and subtle, and they are expressed and perceived differently. We do not claim that our datasets capture the true emotions of the speakers, fully represent language use across the 32 languages, or cover all possible emotions. We discuss the ethical considerations extensively in Section 9.